1) What charge is stored in a 180 µF180 µF capacitor when 120 V is applied to it?

2) Calculate the voltage applied to a 2.00 µF capacitator when it holds 3.10 µC of charge

3) Find the capacitance of a parallel plate capacitor having plates of area 5.00 m^2 that are separated by 0.100 mm of Teflon.

1. What charge is stored in a 180uF capacitor when 120V is applied to it?

2. Calculate the voltage applied to a 2.00 uF capacitor when it holds 3.10 uC of charge

3. Find the capacitance of a parallel plate capacitor having plates of area 5.00 m^2 that are separated by 0.100 mm of Teflon.
Answer #1

Charge

Q=CV=180*120= 21600= 0.0216 microcolunb

Voltage applied to a 2 microferrad

V=Q/C= 3.10/2= 1.55 volt

Capacitance=

C=E*A/d, E= Er*E°= 2.1*8.854*10^-12

C= 2.1*8.854*10^-12 * 5/0.1*10^-3= 0.93 Microferrad